OpenStudy (anonymous):

Plane flies 405 miles with wind and 315 miles against wind in same length of time. If speed of wind is 20 mph, find speed of plane in still air. Show me how to set up? i can solve it :) Thanks!

OpenStudy (anonymous):

I'm confused, i think its supposed to look like rational expression, my teacher taught it that way, i think

OpenStudy (apoorvk):

let the time taken be 't' and plane's actual velocity be 'v' so effectively, with wind, v+20=405/t and against wind, v-20=315/t now solve the equation for 'v' and 't', and enjoy! (post and verify your answer if you need)

OpenStudy (anonymous):

Ill solve it, brb

OpenStudy (anonymous):

My answer: 160

OpenStudy (anonymous):

that is not what i get time is distance divided by rate so you have to expressions for time \(t=\frac{405}{r+20}=\frac{315}{r-20}\)

OpenStudy (anonymous):